# Reaching Curzon-Ahlborn limit in linear response and Whitney limit in nonlinear response in edge mode quantum thermoelectrics and refrigeration

### v2

@@ Abstract
Figure 2: Parametric plot of $\frac{\eta}{\eta_{c}}$ vs $\frac{P}{P_{max}}$ (a) QH, (b) QSH QHE. the blue (red) dot-dashed curve is for $V=0.057k_{B}\theta/e$, the blue (red) dashed curve for $V=0.081k_{B}\theta/e$ and the blue (red) dotted curve for $V=0.096k_{B}\theta/e$. The blue (red) line shows the bound to the power and efficiency. Parametric plot of $\frac{\eta^{r}}{\eta^{r}_{c}}$ vs $\frac{\textbf{J}}{\textbf{J}|_{\eta^{r}_{max}}}$ for (c) QH and (d) QSH QR, for QH (QSH) QR, blue (red) line is for $V=0.115k_{B}\theta/e$. Parameters taken are $\omega=0.1k_{B}\theta/\hbar$, $\theta_{1}=1.1K,\theta_{2}=1.0K$, $\mu=0$, $E_{1}=E_{2}=k_{B}\theta$.
@@
Figure 3: Parametric plot of $\frac{\eta}{\eta_{c}}$ vs $\frac{P}{P_{max}^{Wh}}$ for (a) QH and (b) QSH, QHE. For QH (QSH) QHE, the blue (red) dot-dashed curve is for $V=1.14k_{B}\theta/e$, the blue (red) dashed curve for $V=1.77k_{B}\theta/e$ and the blue (red) dotted curve for $V=3.7k_{B}\theta/e$. The blue (red) line shows the bound to the power and efficiency. Parametric plot for $\frac{\eta^{r}}{\eta^{r}_{c}}$ vs $\frac{\textbf{J}}{\textbf{J}^{Wh}_{max}}$ for (c) QH and (d) QSH, QR. For QH (QSH) QR, blue (red) line is for $V=20k_{B}\theta/e$. Parameters taken are $\omega=\frac{0.1k_{B}\theta}{h}$, $\theta_{1}=2K,\theta_{2}=1K$, $\mu=0$, $E_{1}=E_{2}=k_{B}\theta$.
Non-linear Transport in QH and QSH Setups: There has been ongoing work in this direction on nonlinear thermoelectric transport [18, 19, 32, 33, 34] in both QH and QSH setups. In a two-terminal QH setup, significant deviations from the Onsager-Casimir relation have been observed [32] with a similar investigation done in QSH setup [33]. In a three-terminal QH setup [34] with a voltage probe, efficient operation as a QR in the weakly non-linear regime has been demonstrated. Additionally, a normal metal Aharonov-Bohm heat engine exhibits enhanced thermoelectric performance in the nonlinear regime [29], surpassing its linear regime counterpart [30].
In the linear response regime, electron transport depends solely on the electron’s kinetic energy. However, in the nonlinear regime, it relies on both the kinetic energy and the interaction potential ($U^{\prime}$) of the sample where the electrons flow. It implies that whereas in the linear regime, the QPC energy level $E_{l}$ shifts via the gate voltage: $E_{l}\longrightarrow E_{l}+eV_{g}$, in the nonlinear regime, the interaction potential $U^{\prime}$ shifts the QPC energy level $E_{l}$ as $E_{l}\longrightarrow E_{l}+eU^{\prime}$, with $U^{\prime}=U+u_{g}V_{g}$, where $u_{g}$ is the characteristic potential determining the response of $U^{\prime}$ to the gate voltage $V_{g}$. For the form of $U$, see Eq. (54) of SM. Introducing a gate voltage helps nullify the interaction potential ($U$) in the nonlinear regime, as explained in Sec. II.2.2 of SM and Ref. [29]. From discussion in Sec. II.2.2 of SM , the gate voltage must be of the form $V_{g}=(-U+V_{g}^{\prime})/u_{g}$, where the first term inside the parentheses nullifies the interaction potential $U$, while the second term raises the energy level $E_{l}$: $E_{l}\longrightarrow E_{l}+eV_{g}^{\prime}$ for the nonlinear response regime. We have given an estimation for the interaction potentials in two-terminal QH and three-terminal QH setups with a voltage-temperature probe in SM (see, Tables I and II of SM) for different applied biases. Similarly, we provide an estimate of the interaction potential for two-terminal QSH and three-terminal QSH setups with voltage-temperature probe in SM (Tables III and IV of SM).
@@
In QH setups, as shown in Refs. [18, 19], there exists a bound to the maximum power ($P_{max}=P_{max}^{Wh}$), given by $0.0642\pi^{2}Nk_{B}^{2}(\theta_{1}-\theta_{2})^{2}/h$. Similarly, the bound to the efficiency at maximum power ($\eta|_{P_{max}}$) is found at $P_{max}^{Wh}$, yielding $\eta|{P_{max}^{Wh}}=\eta_{c}/(1+0.936(1+\theta_{2}/\theta_{1}))$, where $\theta_{1}$ and $\theta_{2}$ are the temperatures of the hot and cold reservoirs, and $N$ is the number of edge modes. In our case, with $\theta_{1}=2K$, $\theta_{1}=1K$, and $N=1$, we have $P_{max}^{Wh}=0.632k_{B}^{2}/h$ and $\eta|_{P_{max}^{Wh}}=0.41\eta_{c}$. For power output $P\ll P_{max}^{Wh}$, the bound to $\eta_{max}$ is given by $\eta|_{max}^{Wh}=\eta_{c}\left(1-0.478\sqrt{\theta_{2}P/\theta_{1}P_{max}^{Wh}}\right)$. These quantities were derived for a two-terminal QH setup with a constriction and boxcar-type transmission (see, Sec. II.2.1 of SM and also Refs. [18, 19]). We extend the same approach to a two-terminal QSH setup and prove that the bounds to $P_{max}$, $\eta|_{P_{max}}$, and $\eta_{max}$ are exactly similar to the QH case (see, Sec. II.2.1 of SM).
@@
Adjusting the gate voltage to nullify the interaction potential shifts the energy level of the QPC-like constriction. This is explained in Sec. II.2.2 of SM, initially demonstrated in a two-terminal nonlinear QH setup and extended to a three-terminal configuration (Sec. II.2.4 of SM). A similar approach is applied to a two-terminal QSH setup in Sec. III.2.2 of SM, later extended to a three-terminal configuration (Sec. III.2.4 of SM). We examine $P_{max}$, $\eta|_{P_{max}}$ and $\eta_{max}$ numerically. In Fig. 3(a), for the QH setup, maximum power approaches the Whitney limit at a specific voltage bias $V=1.14k_{B}\theta/e$ (blue dot-dashed curve). With increasing voltage bias $V$, power decreases and efficiency increases (see blue dashed line for $V=1.77k_{B}\theta/e$ and blue dotted for $V=3.7k_{B}\theta/e$ in Fig. 3(a)). The QHE is more efficient as $\eta|_{P_{max}}$ and $\eta_{max}$ reach $0.41\eta_{c}$ and $0.93\eta_{c}$ respectively. Similarly, in Fig. 3(c), $J_{max}$ reaches $J_{max}^{Wh}$, while the $\eta_{max}^{r}$ is approximately 0.86 $\eta_{c}^{r}$. Details for generating the plots in Fig. 3(a) and (c) are provided in Sec. II.2.4 of SM.
@@
Experimental Realization and Conclusion: In three-terminal QH and QSH setups with QPC-like constrictions, we achieve optimal thermoelectric performance approaching efficiency benchmarks for QHE and coefficient of performance for QR, both in linear (Fig. 2) and nonlinear (Fig. 3) response regimes. For nonlinear response, maximum power generated approaches Whitney bound $P_{max}^{Wh}$ (see Fig. 3(a), Table 1) and maximum cooling power approaches $J_{max}^{Wh}$ (see Fig. 3(c), Table 1) for QH setup. In QSH, maximum power as QHE decreases from $P_{max}^{Wh}$, yet efficiency at maximum power remains $0.41\eta_{c}$ (Curzon-Ahlborn efficiency equivalent) (see Fig. 3(b), Table 1). Similarly, as QR, maximum cooling power is below $J_{max}^{Wh}$ (Fig. 3(d), Table 1). This power suppression (both $P_{max}^{Wh}$ and $J_{max}^{Wh}$) arises as a result of helical edge mode transport with a VT probe as there is a finite spin current in three-terminal QSH setup, which is absent in chiral case. Designing a VT probe experimentally is slightly challenging due to potential heat flow between reservoir and surroundings, altering its temperature [16]. We employ QPC-type tunneling, utilized in prior works [22, 23, 24, 25] to study electron transport in 2DEGs, and it’s easier to implement experimentally and is controlled electrostatically with split gates [36, 37, 38].
In the linear response regime, an optimal quantum thermoelectric material is characterized by a figure of merit ($Z\theta$) significantly greater than 1. Theoretical proposals report $Z\theta$ values of 4 in ferromagnet-superconductor junctions (Ref. [39]) and up to 40-50 in other studies (Refs. [40], [41]). Materials with $Z\theta\gg 1$ are highly desirable for optimal thermoelectric performance as both quantum heat engines and refrigerators. To summarize, the main take home messages of this letter are: 1. Use of QPC instead of boxcar type transmission [38, 24, 22, 25], which is much easier to implement, 2. Investigating thermoelectric transport in the non-linear response regime for QSH setups, which has never been attempted, 3. Investigation of QH and QSH setups as QHE and QR with a voltage-temperature probe in the non-linear response regime, which has again never been explored before and finally 4. An estimate of interaction potential and the applied gate voltage to nullify it in 2T QSH setup (see, Sec. III.2.2 of SM ) and 3T QH (see, Sec. II.2.4 of SM ) and QSH setups (see, Sec. III.2.4 of SM ) with VTP for the first time.
@@
In conclusion, this letter proposes achieving the Curzon-Ahlborn (CA) limit in linear and the corresponding Whitney bound in the nonlinear response regimes through a setup applicable to three-terminal QH and QSH systems, extendable to multiterminal setups with multiple VT probes. Furthermore, this approach can be applied to various edge mode systems, including trivial helical [44, 45, 46] and anomalous modes [47, 48, 49].

### v3

@@ Abstract
Figure 3: Parametric plot of $\frac{\eta}{\eta_{c}}$ vs $\frac{P}{P_{max}}$ (a) QH, (b) QSH QHE. the blue (red) dot-dashed curve is for $V=0.057k_{B}T/e$, the blue (red) dashed curve for $V=0.081k_{B}T/e$ and the blue (red) dotted curve for $V=0.096k_{B}T/e$. The blue (red) line shows the bound to the power and efficiency. Parametric plot of $\frac{\eta^{r}}{\eta^{r}_{c}}$ vs $\frac{\textbf{J}}{\textbf{J}|_{\eta^{r}_{max}}}$ for (c) QH and (d) QSH QR, for QH (QSH) QR, blue (red) line is for $V=0.115k_{B}T/e$. Parameters taken are $\omega=0.1k_{B}T/\hbar$, $T_{1}=1.1K,T_{2}=1.0K$, $\mu=0$, $E_{1}=E_{2}=k_{B}T$.
@@
Figure 4: Parametric plot of $\frac{\eta}{\eta_{c}}$ vs $\frac{P}{P_{max}^{Wh}}$ for (a) QH and (b) QSH, QHE. For QH (QSH) QHE, the blue (red) dot-dashed curve is for $V=1.14k_{B}T/e$, the blue (red) dashed curve for $V=1.77k_{B}T/e$ and the blue (red) dotted curve for $V=3.7k_{B}T/e$. The blue (red) line shows the bound to the power and efficiency. Parametric plot for $\frac{\eta^{r}}{\eta^{r}_{c}}$ vs $\frac{\textbf{J}}{\textbf{J}^{Wh}_{max}}$ for (c) QH and (d) QSH, QR. For QH (QSH) QR, blue (red) line is for $V=20k_{B}T/e$. Parameters taken are $\omega=\frac{0.1k_{B}T}{h}$, $T_{1}=2K,T_{2}=1K$, $\mu=0$, $E_{1}=E_{2}=k_{B}T$.
Non-linear Transport in QH and QSH Setups: There has been ongoing work in this direction on nonlinear thermoelectric transport [13, 14, 50, 51, 52, 53, 54, 55, 56, 57] in both QH and QSH setups. In a two-terminal QH setup, significant deviations from the Onsager-Casimir relation have been observed [50] with a similar investigation done in QSH setup [51]. In a three-terminal QH setup [52] with a voltage probe, efficient operation as a QR in the weakly non-linear regime has been demonstrated. Additionally, a normal metal Aharonov-Bohm heat engine exhibits enhanced thermoelectric performance in the nonlinear regime [11], surpassing its linear regime counterpart [12].
In the linear response regime, electron transport depends solely on the electron’s kinetic energy. However, in the nonlinear regime, it relies on both the kinetic energy and the interaction potential ($U^{\prime}$) of the sample where the electrons flow [53, 54, 55, 56, 57, 58]. It implies that whereas in the linear regime, the QPC energy level $E_{l}$ shifts via the gate voltage: $E_{l}\longrightarrow E_{l}+eV_{g}$, in the nonlinear regime, the interaction potential $U^{\prime}$ shifts the QPC energy level $E_{l}$ as $E_{l}\longrightarrow E_{l}+eU$, with $U=U^{\prime}+u_{g}V_{g}$, where $u_{g}$ is the characteristic potential determining the response of $U$ to the gate voltage $V_{g}$. For the form of $U$, see Eq. (54) of SM. Introducing a gate voltage helps nullify the interaction potential ($U^{\prime}$) in the nonlinear regime, as explained in Sec. I.2.2 of SM and Ref. [11]. From discussion in Sec. I.2.2 of SM, the gate voltage must be of the form $V_{g}=(-U^{\prime}+V_{g}^{\prime})/u_{g}$, where the first term inside the parentheses nullifies the interaction potential $U^{\prime}$, while the second term raises the energy level $E_{l}$: $E_{l}\longrightarrow E_{l}+eV_{g}^{\prime}$ for the nonlinear response regime. We have given an estimation for the interaction potentials in two-terminal QH and three-terminal QH setups with a voltage-temperature probe in SM (see, Tables 2 and 3 of SM for different applied biases. Similarly, we provide an estimate of the interaction potential for two-terminal QSH and three-terminal QSH setups with voltage-temperature probe in SM (Tables 4 and 5 of SM.
@@
In QH setups, as shown in Refs. [13, 14], there exists a bound to the maximum power ($P_{max}=P_{max}^{Wh}$), given by $0.0642\pi^{2}Nk_{B}^{2}(T_{1}-T_{2})^{2}/h$. Similarly, the bound to the efficiency at maximum power ($\eta|_{P_{max}}$) is found at $P_{max}^{Wh}$, yielding $\eta|{P_{max}^{Wh}}=\eta_{c}/(1+0.936(1+T_{2}/T_{1}))$, where $T_{1}$ and $T_{2}$ are the temperatures of the hot and cold reservoirs, and $N$ is the number of edge modes. In our case, with $T_{1}=2K$, $T_{1}=1K$, and $N=1$, we have $P_{max}^{Wh}=0.632k_{B}^{2}/h$ and $\eta|_{P_{max}^{Wh}}=0.41\eta_{c}$. For power output $P\ll P_{max}^{Wh}$, the bound to $\eta_{max}$ is given by $\eta|_{max}^{Wh}=\eta_{c}\left(1-0.478\sqrt{T_{2}P/T_{1}P_{max}^{Wh}}\right)$. These quantities were derived for a two-terminal QH setup with a constriction and boxcar-type transmission (see, Sec. I.2.1 of SM and also Refs. [13, 14]). We extend the same approach to a two-terminal QSH setup and prove that the bounds to $P_{max}$, $\eta|_{P_{max}}$, and $\eta_{max}$ are exactly similar to the QH case (see, Sec. I.2.1 of SM).
@@
Adjusting the gate voltage to nullify the interaction potential shifts the energy level of the QPC-like constriction. This is explained in Sec. I.2.2 of SM, initially demonstrated in a two-terminal nonlinear QH setup and extended to a three-terminal configuration (Sec. I.2.4 of SM. A similar approach is applied to a two-terminal QSH setup in Sec. II.2.2 of SM, later extended to a three-terminal configuration (Sec. II.2.4 of SM. We examine $P_{max}$, $\eta|_{P_{max}}$ and $\eta_{max}$ numerically. In Fig. 4(a), for the QH setup, maximum power approaches the Whitney limit at a specific voltage bias $V=1.14k_{B}T/e$ (blue dot-dashed curve). With increasing voltage bias $V$, power decreases and efficiency increases (see blue dashed line for $V=1.77k_{B}T/e$ and blue dotted for $V=3.7k_{B}T/e$ in Fig. 4(a)). The QHE is more efficient as $\eta|_{P_{max}}$ and $\eta_{max}$ reach $0.41\eta_{c}$ and $0.93\eta_{c}$ respectively. Similarly, in Fig. 4(c), $J_{max}$ reaches $J_{max}^{Wh}$, while the $\eta_{max}^{r}$ is approximately 0.86 $\eta_{c}^{r}$. Details for generating the plots in Fig. 4(a) and (c) are provided in Sec. I.2.4 of S/M.
@@
Experimental Realization and Conclusion: In three-terminal QH and QSH setups with QPC-like constrictions, we achieve optimal thermoelectric performance approaching efficiency benchmarks for QHE and coefficient of performance for QR, both in linear (Fig. 3) and nonlinear (Fig. 4) response regimes. For nonlinear response, maximum power generated approaches Whitney bound $P_{max}^{Wh}$ (see Fig. 4(a), Table 1) and maximum cooling power approaches $J_{max}^{Wh}$ (see Fig. 4(c), Table 1) for QH setup. In QSH, maximum power as QHE decreases from $P_{max}^{Wh}$, yet efficiency at maximum power remains $0.41\eta_{c}$ (Van den Broeck efficiency equivalent) (see Fig. 4(b), Table 1). Similarly, as QR, maximum cooling power is below $J_{max}^{Wh}$ (Fig. 4(d), Table 1). This power suppression (both $P_{max}^{Wh}$ and $J_{max}^{Wh}$) arises as a result of helical edge mode transport with a VT probe as there is a finite spin current in three-terminal QSH setup, which is absent in chiral case. Designing a VT probe experimentally is slightly challenging due to potential heat flow between reservoir and surroundings, altering its temperature [32].
In the linear response regime, an optimal quantum thermoelectric material is characterized by a figure of merit ($ZT$) significantly greater than 1. Theoretical proposals report $ZT$ values of 4 in ferromagnet-superconductor junctions (Ref. [19]) and up to 40-50 in other studies (Refs. [20], [21]). Materials with $ZT\gg 1$ are highly desirable for optimal thermoelectric performance as both quantum heat engines and refrigerators. To summarize, the main take home messages of this letter are: 1. Use of QPC instead of boxcar type transmission [48, 39, 37, 40], which is much easier to implement, 2. Investigating thermoelectric transport in the non-linear response regime for QSH setups, which has never been attempted, 3. Investigation of QH and QSH setups as QHE and QR with a voltage-temperature probe in the non-linear response regime, which has again never been explored before and finally 4. An estimate of interaction potential and the applied gate voltage to nullify it in 2T QSH setup (see, Sec. II.2.2 of SMand 3T QH (see, Sec. I.2.4 of SM and QSH setups (see, Sec. II.2.4 of SM with VTP for the first time.
@@
In conclusion, this letter proposes achieving the Van den Broeck (VDB) limit in linear and the corresponding Whitney bound in the nonlinear response regimes through a setup applicable to three-terminal QH and QSH systems, extendable to multiterminal setups with multiple VT probes. Furthermore, this approach can be applied to various edge mode systems, including trivial helical [62, 63, 64] and anomalous modes [65, 66, 67].